OpenStudy (anonymous):

if sintheta=3/4 than what are the values of cos theta?

OpenStudy (anonymous):

@campbell_st can you help??

OpenStudy (campbell_st):

ok.... really simple draw a triangle  the concerning bits are values.... cos and sin are both positive in the 1st quadrant and in the 2nd quadrant sin is positive and cos is negative... so you will have an answer cos = a/4 and cos = -a/4 hope it helps

OpenStudy (anonymous):

thank you so much!